Find the product of two consecutive positive even integers is 288. Find the integers.
Answer by Alan3354(30993)

Find the product of two consecutive positive even integers is 288. Find the integers.
----------------
There are 2 ways to do this.
Method #1
Find the square root of 288. It's just under 17 (17^ = 289). It's the even numbers just under and over 17, 16 & 18.
Done.
-------------
Method #2
x*(x+2) = 288
x^2 + 2x - 288 - 0
(x-16)*(x + 18) = 0
x = 16 (2nd # is 18)
A lot more work.